# ResQuNNs:Towards Enabling Deep Learning in Quantum Convolution Neural Networks

### v4

@@ Abstract
In this paper, we introduce a framework aimed at enhancing the performance of Quantum Convolutional Neural Networks a.k.a Quanvolutional Neural Networks (QuNNs) by incorporating trainable quanvolutional layers and addressing the significant trainability challenges inherent in multi-layered QuNNs. Current state-of-the-art QuNNs predominantly utilize static quanvolutional layers, which limit their feature extraction capabilities. Our proposed approach mitigates this limitation by enabling the training of these layers, thereby substantially improving the scalability and learning potential of QuNNs. However, deep QuNNs—those with multiple trainable quanvolutional layers—encounter complexities in gradient-based optimization, primarily due to restricted gradient flow across these layers. To resolve this, we propose Residual Quanvolutional Neural Networks (ResQuNNs), leveraging the concept of residual learning, which facilitates the flow of gradients by adding skip connections between these layers. The insertion of residual blocks between quanvolutional layers ensures enhanced gradient access through all the layers of the network, thus enabling deep learning in QuNNs and eventually leading to an improved training performance. Moreover, we provide empirical evidence on the strategic placement of these residual blocks within ResQuNNs. Through extensive experimentation, we identify efficient configurations of residual connections, which enable gradients flow across all the layers in the network that eventually results in efficient training. Our results mark a substantial step forward in the evolution of quantum deep learning, offering new avenues for both theoretical development and practical quantum computing applications.
Index Terms:
@@
I Introduction
Quantum Convolutional Neural Networks (QCNNs) represent an innovative fusion of quantum computing principles with traditional convolutional neural network (CNN) architectures [1, 2]. These hybrid classical-quantum models utilize parameterized quantum circuits (PQCs) a.k.a ansatz, composed of parameterized rotation gates such as $R_{x}$, $R_{y}$, or $R_{z}$, functioning analogously to convolutional and pooling operations in classical CNNs [3, 4]. The optimization, however, is performed on classical computers [5, 6].
A notable advantage of QCNNs lies in their ability to harness the vast Hilbert space offered by quantum mechanics, which surpasses the capabilities of classical CNNs [7]. This feature allows QCNNs to capture spatial relationships in image data more effectively. Furthermore, the potential future inclusion of higher-level quantum systems, such as qutrits or ququads [8, 9], promises even more sophisticated image comprehension capabilities.
There are two primary variants of QCNNs, each leveraging quantum mechanics principles like superposition, entanglement, and interference. The first type, Quantum-Inspired CNNs, proposed in [3], mirrors the structure of classical CNNs but replaces convolutional and pooling layers with deeper PQCs. This design demands higher qubit coherence to fully exploit quantum parallelism and entanglement, posing a challenge for current quantum hardware. Nevertheless, a significant research has been dedicated to explore this very architectural framework for various applications [5, 10]. The second type, Quanvolutional Neural Networks (QuNNs), first introduced in [11], and visually depicted in Fig. 1(a), focuses on quantum-enhanced convolutional layers. These networks encode classical image pixels into quantum states, apply unitary transformations via PQCs, and then measure the qubits to create output image channels, which are subsequently processed by classical fully connected layers. This approach offers a more feasible implementation with current quantum technology. Therefore, in this paper, we also delve into QuNNs with a focus on identifying the challenges faced in training QuNNs and proposing effective solutions to address these issues.
@@ I Introduction
An abstract view of the workflow of quanvolution layer in QuNNs is presented in Fig. 1(b), which is four step process: (1) the original image $U$ is divided into spatially located subsections $u_{x}$. Each subsection is a $d\times d$ matrix with $d>1$, (2) these image subsections are then encoded into a quantum state ($q$), e.g., for an encoding function $E$, the quantum state can be defined as $q_{x}=E(u_{x})$. We encode the input features into the qubit rotation angles using $R_{y}(\theta)$. (3) the encoded state $q_{x}$ then undergoes a unitary transformation via a quantum circuit a.k.a ansatz $Q$ which can be described as; $f_{x}=Q(q_{x})=Q(E(u_{x}))$, (4) decoding is then performed via measurements which can be described as, $D(f_{x})=D(Q(E(u_{x})))$. By repeating these steps for different subsections, the full input image is traversed, producing an outcome organized as a multi-channel image. The output of this quanvolution is usually processed by a classical neural network to produce the final output.
This QuNN architecture, widely adopted in state-of-the-art research, is extensively utilized for various applications such as 2D and 3D radiological image classification [12], high energy physics data analysis [13], and object detection and classification [14, 15]. In these contexts, the quantum circuit in QuNNs primarily functions as a feature extractor and is not trained during the quanvolution operation. Consequently, the learning process is primarily driven by the subsequent classical network.
Although there are recent studies that employ trainable quantum circuits in quanvolutional layers, to the best of our knowledge, no existing literature addresses multi-layered QuNNs. Table I some summarizes recent works utilizing QuNNs with both trainable and non-trainable quanvolutional layers. To achieve full scalability in QuNNs, it is crucial to ensure the trainability of underlying quanvolutional layers in multi-layered architectures.
TABLE I: Recent State-of-the-art QuNNs employing trainable and untrainable quanvolution layers
@@
I-A Motivational Case Study
In Fig. 2, we assess the impact of trainable versus untrainable quanvolutional layers in QuNNs. To this end, we train QuNNs with both trainable (allowing weight updates during backpropagation) and untrainable (frozen weights, no updates) quanvolutional layers. This training is performed on $1000$ randomly chosen images from MNIST dataset over $30$ iterations, using an $80\%-20\%$ training-validation split. The Adam optimizer with a $0.01$ learning rate was employed. We utilized a $2\times 2$ kernel resulting in $4$-qubit quanvolutional layers, with each layer comprising of $4$ parameterized $R_{y}$ gates and $3$ $CNOT$ gates for (nearest neighbor) qubit entanglement. The overall depth of quantum circuit in the quanvolutional layer is set to $2$, accounting to a total of $8$ single and $6$ two-qubit gates.
@@ I-A Motivational Case Study
While a single trainable quanvolutional layer enhances QuNNs’ training efficacy, as depicted in Fig. 2, scalability of QuNNs requires multiple trainable layers. However, challenges emerge when multiple quanvolutional layers are trained concurrently, as gradients are exclusively accessible only in the last quanvolutional layer, thus limiting the optimization process only to this layer. Fig. 3 demonstrates this by showing gradients for two and three trainable quanvolutional layers.
We observe that the gradients of only the final quanvolutional layer are accessible, while those of the preceding layers are effectively None. The restricted gradient flow means that in multi-layered QuNNs, only the last quanvolutional layer contributes in the learning process, leaving earlier layers remain uninvolved in optimization. This presents a significant challenge in effectively training multi-layered QuNNs, thereby necessitating innovative approaches to address this challenge for scalable and effective QuNN architectures.
@@
I-B Novel Contributions
Our primary contributions encompass the introduction of trainable quanvolutional layers, identification of challenges with multiple quanvolutional layers, the development of Residual Quanvolutional Neural Networks (ResQuNNs) to address these challenges, and empirical findings on the strategic placement of residual blocks to further enhance training performance. An overview of our contributions is shown in Fig. 4.
  • Trainable Quanvolutional Layers: We first make the quanvolutional layers trainable in QuNNs, addressing the challenge of enhancing their adaptability within QuNNs and improving their training performance.
  • Challenges with Multiple Quanvolutional Layers: We identify a key challenge in optimizing multiple quanvolutional layers, highlighting the complexities associated with accessing gradients across these layers. In QuNN architectures with multiple trainable quanvolutional layers, the gradients do not flow through all the layers of the network, making deeper QuNNs unsuitable for gradient-based optimization techniques.
  • Residual Quantum Convolutional Neural Networks (ResQuNN): To overcome the gradient accessibility issue, this paper introduces Residual Quantum Convolutional Neural Networks (ResQuNN), drawing inspiration from classical residual neural networks.
  • Residual Blocks for Gradient Accessibility: In ResQuNNs, the residual blocks between quanvolutional layers are incorporated to facilitate comprehensive gradient access throughout the network, thereby enabling the deep learning in QuNNs and eventually improving training performance of QuNNs. These residual blocks combine the output of existing quanvolutional layers with their own or previous layer’s input before forwarding it to the subsequent quanvolutional layer.
  • Strategic Placement of Residual Blocks: We conducted extensive experiments to determine appropriate locations for inserting residual blocks in ResQuNNs which ensure comprehensive gradient throughout the network.
  • Both Classical and Quantum Postprocessing: State-of-the-art QuNNs utilize classical networks to postprocess the output of quanvolutional layers. However, to demonstrate the effectiveness of the residual approach in multi-layered QuNNs (with no influence on the overall learning from classical layers), other than classical postprocessing layer, we employ the quantum circuit itself for postprocessing the results of the quanvolutional layers.
II Related Work
@@
III ResQuNN Methodology
The detailed overview of ResQuNN’s methodology is presented in Fig. 5. Below we discuss different steps of our methodology in detail.
III-A Trainable Quanvolutional Layers in QuNNs.
The first step is to make the quanvolutional layers trainable. The majority of works in the literature utilizes these layers for efficient feature extraction and pass them to classical layers, however, the quanvolution operation itself is not trainable. In order to enhance the adaptability of these layers within QuNNs it is important to make these layers trainable to harness their full potential. Making the quanvolutional layer trainable entails making the quantum circuit (labeled as 1 in Fig. 5) trainable. We do this by making the quanvolution layer as a pytorch layer which is later used as a standard layer during the model construction. The trainable QuNN utilized in this study is depicted in purple shaded region of Fig. 5, and operates through the following steps.
@@ III ResQuNN Methodology
    $U_{\text{circuit}_{1}}=\left(\bigotimes_{i=1}^{3}CNOT_{i,i+1}\right)\cdot\left(\bigotimes_{i=1}^{4}RX(\theta_{i})\right)\cdot\left(\bigotimes_{i=1}^{4}RY(x_{i})\right)$
@@
    where $l$ represents the total number of layers, i.e., number of times the circuit is repeated before measurement. The size of output after every quanvloutional layers is detailed in Table II. |$\ket{\psi_{\text{QCL1}}}$ is the final state state after the first quanvolutionla layer before measurement and upon measurement the quantum state collapses and produce a classical result:
@@
    Classical layer to postprocess quanvolutional layer’s output
    First, we use classical layer to postprocess the output of quanvolutional layer, as highlighted in label 2 in Fig. 5). This is a major practice in state-of-the-art QuNN architectures, where a fully connected neuron layer is used to take the output of the quanvolutional layer and produces the final classification output. The number of neurons in this layer are equal to the total number of classes in the dataset used, which in our case is $10$ since we used MNIST handwritten digit dataset which has $10$ classes (more details in Section IV).
    Let’s assume the output of the quanvolutional layer is a set of measurements $\{M=m_{1},m_{2}\ldots m_{n}\}$, where each $m_{i}$ represents the measurement of a qubit in terms of classical information, such as the expectation values in computational basis. This result is transformed to a vector format and then passed to the fully connected layer with $10$ neurons. The output feature size from quanvolutional layers is different and depends on the kernel size and stride similar to that of classical convolutional layers. We experiment with a kernel size of $2\times 2$ and stride of $2$. The output feature size after each quanvolutional layer in case of different residual configuration (label 5 in Fig. 5) is presented in the Table II.
    TABLE II: Output Feature Sizes for Two Quanvolutional Layer
@@
    Quantum Circuit to postprocess quanvolutional layer’s output
    We also use quantum circuit to postprocess the result from quanvolutional layer, as highlighted by label 3 in Fig. 5. This is primarily for two reasons: firstly to demonstrate the working of fully quantum QuNNs, and secondly, to better understand the importance of residual connections by removing the influence of classical layers being used at the end.
@@
    where $x_{i}$ is the $i^{th}$ element of $x$ and |$\ket{i}$ is the $i^{th}$ computational basis state. We used amplitude encoding for postprocessing the results of quanvolutional layers primarily because it allows to encode $2^{n}$ features in $n$ qubits. Based on Table II, the smallest output feature size from quanvolutional layer(s) is $7\times 7\times 1=49$, which requires atleast $49$ qubits via angle encoding (one qubit per feature) and only $6$ qubits via amplitude encoding. Similarly, the maximum output size of quanvolutional layer is $28\times 28\times 1=784$, requiring $784$ qubits, if encoded in qubit rotation angles, whereas Amplitude Encoding reduces the requirement to merely $10$ qubits. The use of $784$ qubits is not practical in current NISQ devices, therefore we use amplitude encoding to encode the qunavolutional layer’s result for further processing by a quantum circuit. To ensure a consistent impact of the quantum circuit’s post-processing across all residual configurations, a $10$-qubit circuit is utilized for processing the outputs of the quanvolutional layer, corresponding to the maximum feature size of $784$. However, as mentioned earlier, we use $4$ classes, therefore, only $4$ qubits are measured each corresponding to one class. The output of quanvolutional, once encoded, is subjected to the unitary evolution through a simple quantum circuit, which can be described by the following Equation:
@@
III-B Proposed Residual Quanvolutional Neural Networks (ResQuNNs).
We observe that optimizing multiple (two or more) quanvolutional layers presents a significant challenges. One of the main difficulties lies in gradients accessibility across these layers during training, which are essential in gradient-driven optimization techniques. We typically notice that when we train two or more quanvolutional layers, the gradients of only the last layer are accessible. We have shown these results as a primary motivation of this paper in Section I-A.
To overcome the gradient accessibility issue encountered in multi-layered QuNNs, we introduce ResQuNNs. Drawing inspiration from classical residual neural networks, ResQuNNs incorporate residual blocks between quanvolutional layers. The residual blocks combine the output of existing quanvolutional layers with their own or the previous layer’s input before forwarding it to the subsequent quanvolutional layer, as shown in label 4 in Fig. 5. These residual connections play a pivotal role in facilitating comprehensive gradient access and, in turn, improving the overall training performance of QuNNs. The integration of residual blocks ensures that gradients flow efficiently across the network, thereby overcoming the challenges associated with training multiple quanvolutional layers.
III-C Strategic Placement of Residual Blocks.
Within the proposed ResQuNN architecture, not every random residual configuration results in comprehensive gradient flow throughout the network, and the residual blocks needs to be strategically placed between quanvolutional layers. In our empirical investigation, we conduct extensive experimentation to determine the appropriate locations for inserting residual blocks within ResQuNNs comprising two quanvolutional layers, as shown in Fig. 5, labeled 5. This analysis helps us uncover the most effective residual configurations for ResQuNNs, enabling the gradients to flow throughout the proposed ResQuNN architecture, resulting in a deeper architecture and eventually the enhanced training performance. Based on the results of two quanvolutional layers, only the most effective configurations, i.e., those facilitating gradient flow through both layers were chosen for subsequent examination within three-layer quanvolutional frameworks, as denoted by label 6 in Fig. 5.
We then trained ResQuNNs with two quanvolutional layers for all residual configurations for a multiclass classification problem (more on dataset in Section IV) to assess the impact of gradient accessibility, both throughout the entire network (all quanvolutional layers) and within segments of the network (only the last quanvolutional layer), on the learning efficacy of QuNNs.
TABLE III: Dataset and Hyperparemeters Specifications. BS = batch size, Opt = optimizer and LR = learning rate
@@
IV Experimental Setup
An overview of the experimental toolflow utilized in this paper is depicted in Fig. 6. We trained the proposed ResQuNNs on a selected subset of the MNIST dataset. Table III summarizes the dataset and hyperparameters specification used in this paper. The ResQuNNs are trained for $30$ epochs. Adam optimizer with learning rate of $0.01$ is used for the optimization. The objective function chosen for training is cross entropy loss, which can be described (for C classes) by the following equation:

### v5

@@ Abstract
This paper focuses on methidenhances Quantum Convolutional Neural Networks (QuNNs) by introducing trainable quanvolutional layers and addressing the challenges in training deep QuNNs. Traditional QuNNs rely on static quanvolutional layers, limiting their feature extraction capabilities. Our approach enables the training of these layers, significantly improving the scalability and learning potential of QuNNs. However, deep QuNNs face difficulties in gradient-based optimization due to limited gradient flow across multiple layers. To overcome this, we propose Residual Quanvolutional Neural Networks (ResQuNNs), which utilize residual learning by adding skip connections between quanvolutional layers. These residual blocks enhance gradient flow throughout the network, facilitating effective training in deep QuNNs. Moreover, We provide empirical evidence on the optimal placement of these residual blocks, demonstrating how strategic configurations improve gradient flow and lead to more efficient training. Our findings represent a significant advancement in quantum deep learning, opening new possibilities for both theoretical exploration and practical quantum computing applications.
keywords:
@@
1 Introduction
Quantum Convolutional Neural Networks (QCNNs) represent an innovative fusion of quantum computing principles with traditional convolutional neural network (CNN) architectures [1, 2]. These hybrid classical-quantum models utilize parameterized quantum circuits (PQCs) a.k.a ansatz, composed of parameterized rotation gates such as $R_{x}$, $R_{y}$, or $R_{z}$, functioning analogously to convolutional and pooling operations in classical CNNs [3, 4]. The optimization, however, is performed on classical computers [5, 6]. A notable advantage of QCNNs lies in their ability to harness the vast Hilbert space offered by quantum mechanics, which surpasses the capabilities of classical CNNs [7]. This feature allows QCNNs to capture spatial relationships in image data more effectively. Furthermore, the potential future inclusion of higher-level quantum systems, such as qutrits or ququads [8, 9], promises even more sophisticated image comprehension capabilities.
There are two primary variants of QCNNs, each leveraging quantum mechanics principles like superposition, entanglement, and interference. The first type, Quantum-Inspired CNNs, proposed in [3], mirrors the structure of classical CNNs but replaces convolutional and pooling layers with deeper PQCs. This design demands higher qubit coherence to fully exploit quantum parallelism and entanglement, posing a challenge for current quantum hardware. Nevertheless, a significant research has been dedicated to explore this very architectural framework for various applications [5, 10]. The second type, Quanvolutional Neural Networks (QuNNs), first introduced in [11], and visually depicted in Figure 1(a), focuses on quantum-enhanced convolutional layers. These networks encode classical image pixels into quantum states, apply unitary transformations via PQCs, and then measure the qubits to create output image channels, which are subsequently processed by classical fully connected layers. This approach offers a more feasible implementation with current quantum technology. Therefore, in this paper, we also delve into QuNNs with a focus on identifying the challenges faced in training QuNNs and proposing effective solutions to address these issues.
@@ 1 Introduction
An abstract view of the workflow of quanvolution layer in QuNNs is presented in Figure 1(b), which is four step process: (1) the original image $U$ is divided into spatially located subsections $u_{x}$. Each subsection is a $d\times d$ matrix with $d>1$, (2) these image subsections are then encoded into a quantum state ($q$), e.g., for an encoding function $E$, the quantum state can be defined as $q_{x}=E(u_{x})$. We encode the input features into the qubit rotation angles using $R_{y}(\theta)$. (3) the encoded state $q_{x}$ then undergoes a unitary transformation via a quantum circuit a.k.a ansatz $Q$ which can be described as; $f_{x}=Q(q_{x})=Q(E(u_{x}))$, (4) decoding is then performed via measurements which can be described as, $D(f_{x})=D(Q(E(u_{x})))$. By repeating these steps for different subsections, the full input image is traversed, producing an outcome organized as a multi-channel image. The output of this quanvolution is usually processed by a classical neural network to produce the final output.
The QuNN architecture depicted in Figure 1, widely adopted in state-of-the-art research, is extensively utilized, for various applications such as 2D and 3D radiological image classification [12], high energy physics data analysis [13], and object detection and classification [14, 15]. In these contexts, the quantum circuit in QuNNs primarily functions as a feature extractor and is not trained during the quanvolution operation. Consequently, the learning process is primarily driven by the subsequent classical network. Although there are recent studies that employ trainable quantum circuits in quanvolutional layers, to the best of our knowledge, no existing literature addresses multi-layered QuNNs. Table 1 some summarizes recent works utilizing QuNNs with both trainable and non-trainable quanvolutional layers. To achieve full scalability in QuNNs, it is crucial to ensure the trainability of underlying quanvolutional layers in multi-layered architectures.
Table 1: Recent State-of-the-art QuNNs employing trainable and untrainable quanvolution layers
@@
1.1 Motivational Case Study
In Figure 2, we assess the impact of trainable versus untrainable quanvolutional layers in QuNNs. To this end, we train QuNNs with both trainable (allowing weight updates during backpropagation) and untrainable (frozen weights, no updates) quanvolutional layers. This training is performed on $1000$ randomly chosen images from MNIST dataset over $30$ iterations, using an $80\%-20\%$ training-validation split. The Adam optimizer with a $0.01$ learning rate was employed. We utilized a $2\times 2$ kernel resulting in $4$-qubit quanvolutional layers, with each layer comprising of $4$ parameterized $R_{y}$ gates and $3$ $CNOT$ gates for (nearest neighbor) qubit entanglement. The overall depth of quantum circuit in the quanvolutional layer is set to $2$, accounting to a total of $8$ single and $6$ two-qubit gates.
@@ 1.1 Motivational Case Study
While a single trainable quanvolutional layer enhances QuNNs’ training efficacy, as depicted in Figure 2, scalability of QuNNs requires multiple trainable layers. However, challenges emerge when multiple quanvolutional layers are trained concurrently, as gradients are exclusively accessible only in the last quanvolutional layer, thus limiting the optimization process only to this layer. Figure 3 demonstrates this by showing gradients for two and three trainable quanvolutional layers. We observe that the gradients of only the final quanvolutional layer are accessible, while those of the preceding layers are effectively None. The restricted gradient flow means that in multi-layered QuNNs, only the last quanvolutional layer contributes in the learning process, leaving earlier layers remain uninvolved in optimization. This presents a significant challenge in effectively training multi-layered QuNNs, thereby necessitating innovative approaches to address this challenge for scalable and effective QuNN architectures.
1.2 Our Contributions
Our primary contributions encompass the introduction of trainable quanvolutional layers, identification of challenges with multiple quanvolutional layers, the development of Residual Quanvolutional Neural Networks (ResQuNNs) to address these challenges, and empirical findings on the strategic placement of residual blocks to further enhance training performance. An overview of our contributions is shown in Figure 4.
@@ 1.2 Our Contributions
  • Trainable Quanvolutional Layers. We first make the quanvolutional layers trainable in QuNNs, addressing the challenge of enhancing their adaptability within QuNNs and improving their training performance.
  • Challenges with Multiple Quanvolutional Layers. We identify a key challenge in optimizing multiple quanvolutional layers, highlighting the complexities associated with accessing gradients across these layers. In QuNN architectures with multiple trainable quanvolutional layers, the gradients do not flow through all the layers of the network, making deeper QuNNs unsuitable for gradient-based optimization techniques.
  • Residual Quantum Convolutional Neural Networks (ResQuNN). To overcome the gradient accessibility issue, this paper introduces Residual Quantum Convolutional Neural Networks (ResQuNN), drawing inspiration from classical residual neural networks.
  • Residual Blocks for Gradient Accessibility. In ResQuNNs, the residual blocks between quanvolutional layers are incorporated to facilitate comprehensive gradient access throughout the network, thereby enabling the deep learning in QuNNs and eventually improving training performance of QuNNs. These residual blocks combine the output of existing quanvolutional layers with their own or previous layer’s input before forwarding it to the subsequent quanvolutional layer.
  • Strategic Placement of Residual Blocks. We conducted extensive experiments to determine appropriate locations for inserting residual blocks in ResQuNNs which ensure comprehensive gradient throughout the network.
  • Both Classical and Quantum Postprocessing. State-of-the-art QuNNs utilize classical networks to postprocess the output of quanvolutional layers. However, to demonstrate the effectiveness of the residual approach in multi-layered QuNNs (with no influence on the overall learning from classical layers), other than classical postprocessing layer, we employ the quantum circuit itself for postprocessing the results of the quanvolutional layers.
2 Related Work
@@
3 ResQuNN Methodology
The detailed overview of ResQuNN’s methodology is presented in Figure 5. Below we discuss different steps of our methodology in detail.
3.1 Trainable Quanvolutional Layers in QuNNs.
The first step is to make the quanvolutional layers trainable. The majority of works in the literature utilizes these layers for efficient feature extraction and pass them to classical layers, however, the quanvolution operation itself is not trainable. In order to enhance the adaptability of these layers within QuNNs it is important to make these layers trainable to harness their full potential. Making the quanvolutional layer trainable entails making the quantum circuit (labeled as 1 in Figure 5) trainable. We do this by making the quanvolution layer as a pytorch layer which is later used as a standard layer during the model construction. The trainable QuNN utilized in this study is depicted in purple shaded region of Figure 5, and operates through the following steps.
@@ 3 ResQuNN Methodology
    $U_{\text{circuit}_{1}}=\left(\bigotimes_{i=1}^{3}CNOT_{i,i+1}\right)\cdot\left(\bigotimes_{i=1}^{4}RX(\theta_{i})\right)\cdot\left(\bigotimes_{i=1}^{4}RY(x_{i})\right)$ (6)
@@
    where $l$ represents the total number of layers, i.e., number of times the circuit is repeated before measurement. The size of output after every quanvloutional layers is detailed in Table 2. |$\ket{\psi_{\text{QCL1}}}$ is the final state state after the first quanvolutionla layer before measurement and upon measurement the quantum state collapses and produce a classical result:
@@
    3.1.1 Classical layer to postprocess quanvolutional layer’s output
    First, we use classical layer to postprocess the output of quanvolutional layer, as highlighted in label 2 in Figure 5). This is a major practice in state-of-the-art QuNN architectures, where a fully connected neuron layer is used to take the output of the quanvolutional layer and produces the final classification output. The number of neurons in this layer are equal to the total number of classes in the dataset used, which in our case is $10$ since we used MNIST handwritten digit dataset which has $10$ classes (more details in Section 4).
    Let’s assume the output of the quanvolutional layer is a set of measurements $\{M=m_{1},m_{2}\ldots m_{n}\}$, where each $m_{i}$ represents the measurement of a qubit in terms of classical information, such as the expectation values in computational basis. This result is transformed to a vector format and then passed to the fully connected layer with $10$ neurons. The output feature size from quanvolutional layers is different and depends on the kernel size and stride similar to that of classical convolutional layers. We experiment with a kernel size of $2\times 2$ and stride of $2$. The output feature size after each quanvolutional layer in case of different residual configuration (label 5 in Figure 5) is presented in the Table 2.
    Table 2: Output Feature Sizes for Two Quanvolutional Layer
@@ 3.1.1 Classical layer to postprocess quanvolutional layer’s output
    3.1.2 Quantum Circuit to postprocess quanvolutional layer’s output
    We also use quantum circuit to postprocess the result from quanvolutional layer, as highlighted by label 3 in Figure 5. This is primarily for two reasons: firstly to demonstrate the working of fully quantum QuNNs, and secondly, to better understand the importance of residual connections by removing the influence of classical layers being used at the end.
@@ 3.1.2 Quantum Circuit to postprocess quanvolutional layer’s output
    where $x_{i}$ is the $i^{th}$ element of $x$ and |$\ket{i}$ is the $i^{th}$ computational basis state. We used amplitude encoding for postprocessing the results of quanvolutional layers primarily because it allows to encode $2^{n}$ features in $n$ qubits. Based on Table 2, the smallest output feature size from quanvolutional layer(s) is $7\times 7\times 1=49$, which requires atleast $49$ qubits via angle encoding (one qubit per feature) and only $6$ qubits via amplitude encoding. Similarly, the maximum output size of quanvolutional layer is $28\times 28\times 1=784$, requiring $784$ qubits, if encoded in qubit rotation angles, whereas Amplitude Encoding reduces the requirement to merely $10$ qubits. The use of $784$ qubits is not practical in current NISQ devices, therefore we use amplitude encoding to encode the qunavolutional layer’s result for further processing by a quantum circuit. To ensure a consistent impact of the quantum circuit’s post-processing across all residual configurations, a $10$-qubit circuit is utilized for processing the outputs of the quanvolutional layer, corresponding to the maximum feature size of $784$. However, as mentioned earlier, we use $4$ classes, therefore, only $4$ qubits are measured each corresponding to one class. The output of quanvolutional, once encoded, is subjected to the unitary evolution through a simple quantum circuit, which can be described by the following Equation:
@@
3.2 Proposed Residual Quanvolutional Neural Networks (ResQuNNs).
We observe that optimizing multiple (two or more) quanvolutional layers presents a significant challenges. One of the main difficulties lies in gradients accessibility across these layers during training, which are essential in gradient-driven optimization techniques. We typically notice that when we train two or more quanvolutional layers, the gradients of only the last layer are accessible. We have shown these results as a primary motivation of this paper in Section 1.1. To overcome the gradient accessibility issue encountered in multi-layered QuNNs, we introduce ResQuNNs. Drawing inspiration from classical residual neural networks, ResQuNNs incorporate residual blocks between quanvolutional layers. The residual blocks combine the output of existing quanvolutional layers with their own or the previous layer’s input before forwarding it to the subsequent quanvolutional layer, as shown in label 4 in Figure 5. These residual connections play a pivotal role in facilitating comprehensive gradient access and, in turn, improving the overall training performance of QuNNs. The integration of residual blocks ensures that gradients flow efficiently across the network, thereby overcoming the challenges associated with training multiple quanvolutional layers.
3.3 Strategic Placement of Residual Blocks.
Within the proposed ResQuNN architecture, not every random residual configuration results in comprehensive gradient flow throughout the network, and the residual blocks needs to be strategically placed between quanvolutional layers. In our empirical investigation, we conduct extensive experimentation to determine the appropriate locations for inserting residual blocks within ResQuNNs comprising two quanvolutional layers, as shown in Figure 5, labeled 5. This analysis helps us uncover the most effective residual configurations for ResQuNNs, enabling the gradients to flow throughout the proposed ResQuNN architecture, resulting in a deeper architecture and eventually the enhanced training performance. Based on the results of two quanvolutional layers, only the most effective configurations, i.e., those facilitating gradient flow through both layers were chosen for subsequent examination within three-layer quanvolutional frameworks, as denoted by label 6 in Figure 5. We then trained ResQuNNs with two quanvolutional layers for all residual configurations for a multiclass classification problem (more on dataset in Section 4) to assess the impact of gradient accessibility, both throughout the entire network (all quanvolutional layers) and within segments of the network (only the last quanvolutional layer), on the learning efficacy of QuNNs.
Table 3: Dataset and Hyperparemeters Specifications. BS = batch size, Opt = optimizer and LR = learning rate
@@
4 Experimental Setup
An overview of the experimental toolflow utilized in this paper is depicted in Figure 6. We trained the proposed ResQuNNs on a selected subset of the MNIST dataset. Table 3 summarizes the dataset and hyperparameters specification used in this paper. The ResQuNNs are trained for $30$ epochs. Adam optimizer with learning rate of $0.01$ is used for the optimization. The objective function chosen for training is cross entropy loss, which can be described (for C classes) by the following equation: